OpenStudy (anonymous):

use 5 steps for solving. The area of a triangular section of a store cannot be more than 910 feet. if the height of the triangular section is 52 ft. what should the base be?

OpenStudy (anonymous):

use five steps!?!??! A = (1/2)*b*h 910 = (1/2)*b*h 910 = (1/2)*b*52 52*b = 1820 b = 35 ft that's five steps...i guess...

OpenStudy (anonymous):

"due to budget restrictions the perimeter of tammies rectangular garden is limited to 74 yards if the length is one more yard than twice the width what are the demensions of the garden